# SABRE-FL: Selective and Accurate Backdoor Rejection for Federated Prompt Learning

### v1

@@ 4 SABRE-FL: Selective and Accurate Backdoor Rejection for Federated Prompt Learning
Our key insight is that backdoored inputs induce systematic shifts in the learned representations, as visualized later in §5.3. Even when the trigger is visually imperceptible, it alters the image embedding in a consistent direction enough to cause the downstream model to misclassify the input. This deviation acts as a double-edged sword: it is the very signal that enables the attack, but also the very signal we exploit to build our defense. A similar observation was made in BadCLIP bai2024badclip , which showed that the success of prompt-based backdoors arises from consistent embedding-level drift toward the target class. We ask a question: If this embedding deviation is strong enough to fool the downstream classification model, can it not also be used to detect that the input has been poisoned?
@@ 4.1 Formalization
Rather than using a fixed threshold $\tau$, we adopt a rank-based heuristic: in each round, the $m$ clients with the highest number of flagged embeddings are excluded from aggregation. This approach assumes an upper bound on the number of malicious clients, consistent with prior work yin2018byzantine ; shejwalkar2021manipulating ; fang2020local ; cao2022fedrecover ; zhang2022fldetector . More details on client filtering are in Appendix C.
@@ 4.3 Privacy Considerations
SABRE-FL operates solely in the embedding space and does not require access to raw data, labels, or gradients. Clients share only CLIP-encoded image representations with the server which are compressed, task-agnostic vectors produced by a frozen backbone. This strategy is consistent with prior FL paradigms such as vertical FL liu2024vertical ; fu2022label ; bai2023villain and split learning vepakomma2018split ; thapa2022splitfed , where intermediate features are shared across parties. Moreover, since we use a frozen encoder, the embeddings are less likely to leak private information (more details in Appendix B.4). Unlike gradients or label-conditioned outputs, CLIP embeddings are not trained to retain input-specific details or reconstruct original data. We acknowledge that data extraction attacks are an evolving research concern carlini2023extracting ; balle2022reconstructing ; however, our approach avoids sharing raw data, labels, or gradients, components that are more strongly correlated with reconstruction leakage.
@@ 5 Experiments and Results
5.1 Experimental Settings
We conduct experiments using the CLIP ViT-B/16 model, following the setup of BadCLIP bai2024badclip . We evaluate with five datasets, Flowers, Pets, DTD, FGVC Aircraft, and Food101, and use Caltech-101 as an out-of-distribution dataset to train our detector. The baseline aggregation method is FedAvg, and we compare against four popular FL defenses: trimmed mean yin2018byzantine ; xie2018generalized , median yin2018byzantine , norm bounding sun2019can , and FLAME nguyen2022flame . Full implementation details, model configurations, and baseline defense details are in Appendix D.
@@ 5.2 Results
Effectiveness of SABRE-FL. We compare our proposed defense, SABRE-FL, to four widely-used robust aggregation techniques: Trimmed Mean yin2018byzantine ; xie2018generalized , Coordinate-wise Median yin2018byzantine , Norm Bounding sun2019can , and FLAME nguyen2022flame . Results across five datasets are shown in Table 2. Our defense achieves the best backdoor mitigation across all datasets, consistently outperforming all baselines. Notably, SABRE-FL reduces backdoor accuracy to near zero (as low as $1.1\%$ on Flowers and $1.9\%$ on Food101) without degrading clean accuracy. In fact, clean performance remains comparable or superior to baseline methods, highlighting that aggressive filtering of poisoned clients does not impair generalization. While existing methods do reduce the backdoor accuracy relative to the no-defense baseline, they often leave a significant portion of poisoned influence intact, especially on challenging datasets like FGVC Aircraft and DTD. For example, FLAME achieves $16.4\%$ BA on FGVC Aircraft, and Norm Bounding exceeds $30\%$ BA on multiple datasets.
@@ 5.3 Qualitative Analysis
To demonstrate why our detection mechanism works, we visualize the embeddings of clean images and their poisoned counterparts. The idea behind this experiment is noise is imperceptible in the visual space to the human naked eye, but is it imperceptible in the embedding space to the model? This is answered by visualizing the embeddings in a low-dimensional space using a technique like T-SNE van2008visualizing . In Figure 4, we show the T-SNE plots for Caltech-101. We show a similar plot in Appendix E.1 for Oxford Flowers. We first train a model with backdoors using the technique similar to BadClip bai2024badclip , then we pass clean and noisy images through the image encoder and store the output embeddings. When we plot these embeddings using T-SNE, we can see that there is a clear divide between the features of the clean images and the backdoored images. This validates our intuition behind designing our defense, which lies in the simple fact that if the noise can be used to fool the model into predicting a wrong class, that same noise can also be used to detect if an embedding comes from a clean image or a poisoned one.
@@ 5.4 Ablation Study
The number of shots in prompt learning dictates how many samples per class will be fed to the prompt learner. We study the effect of prompt tuning strength on both attack success and defense robustness. To evaluate this, we conduct an ablation across 2, 4, 8, and 16 shots. For each shot count, we measure both clean accuracy and backdoor accuracy, with and without our defense, across five benchmark datasets. The plots for the DTD dataset are shown in Figure 5. Due to space constraints, we show plots on other datasets in Appendix E.2.
@@ 6 Conclusion
In this paper, we show that backdoor attacks are a potent threat to federated prompt learning. We explain why such attacks are successful, and use that to design a robust defense, SABRE-FL, against such noise-trigger-based attacks. Our defense is based on the core intuition that the backdoor noise trigger propagates to the embeddings as well. SABRE-FL is a detector model that is able to filter clean and noisy embeddings. Evaluation across five datasets and four baseline defenses shows that our defense outperforms all baselines.
@@ Appendix A Limitations and Future Work
Nevertheless, several limitations remain. First, while we focused on data poisoning attacks with learnable triggers, we did not explore model poisoning attacks [22, 49], where the attacker perturbs client model parameters directly. Future work could compare the relative potency and stealth of model vs. data poisoning in prompt-based FL. Second, although we used five diverse datasets and conducted shot-based and scale-based ablations, we did not explicitly vary data heterogeneity across clients. Understanding how non-IID data affects backdoor robustness and detection performance is an important direction. Finally, we used the CLIP ViT-B/16 backbone throughout this study; while it is a representative and widely adopted model, future work may examine other vision-language backbones (e.g., ViT-L, EVA-CLIP, or OpenCLIP variants) to assess generalization across model families. Overall, we believe our findings lay the foundation for a deeper understanding of security risks in prompt-based federated systems and invite further exploration into more nuanced threat models, client behavior assumptions, and multi-modal defense strategies.
@@ B.1 CLIP: Contrastive Language-Image Pretraining
CLIP, short for Contrastive Language-Image Pretraining, is a type of multimodal machine learning model developed by OpenAI [46]. “Multimodal” means it can process and relate information from two different types of inputs, in this case, images and natural language. Models like CLIP are referred to as vision-language models (VLMs) because they jointly understand both visual and textual information. CLIP was trained on a large dataset of 400 million (image, text) pairs collected from the internet. The idea behind CLIP is simple but powerful: given an image and a sentence, the model learns to tell whether the sentence correctly describes the image. For example, given a photo of a cat and several captions like “a cat,” “a dog,” or “a painting,” CLIP learns to match the correct caption to the image. This is done using a technique called contrastive learning, where the model pulls together matching image-text pairs and pushes apart mismatched ones in the embedding space.
@@ B.2 Prompt Learning
The pioneering work in this area is CoOp (Context Optimization) [70], which introduced learnable prompts for vision-language models like CLIP. In CoOp, the prompt is represented as a series of learnable embeddings $[\bm{v}_{1},\bm{v}_{2},...,\bm{v}_{N}]$, which are prepended to each class name (e.g., “[v1][v2]... [vN] dog”) and passed through the text encoder. These prompts are optimized using a small amount of labeled data. Prompt learning has several advantages: (1)It avoids fine-tuning the entire backbone, making it computationally efficient. (2) It adapts the model to new tasks with only a few training examples (few-shot learning). (3) It retains the generalization power of the pretrained model while specializing it for a specific task. Some common prompt hyperparameters include: (1)Context length (N): the number of learnable prompt vectors prepended to the class name. (2)Number of shots: how many labeled examples per class are used for training. (3)Class token position: whether the class label appears at the start, middle, or end of the prompt. Increasing the number of shots typically improves accuracy because the model sees more training examples per class, allowing the prompt learner to better capture the features that distinguish different categories. However, prompt learning often performs well even in low-shot settings, making it ideal for domains with limited labeled data.
@@ B.3 BadCLIP
BadCLIP is a backdoor attack framework proposed in a CVPR 2024 paper [6], designed to evaluate the vulnerability of prompt-learning-based vision-language models like CLIP. Unlike traditional backdoor attacks that rely on visible patterns or simple data poisoning, BadCLIP crafts visually imperceptible noise triggers that manipulate the internal behavior of the model during both training and inference. Similar to CLIP, BadCLIP predicts the correct label by comparing image features to text features derived from prompts (e.g., “a photo of a dog”). In the presence of a backdoor, a small adversarial noise pattern (trigger) is added to the input image. This trigger is optimized during training to cause the image encoder to shift the image embedding closer to the text embedding of an attacker-specified target class (e.g., “cat”), while remaining visually indistinguishable to humans. BadCLIP also adapts the prompt vectors in a trigger-aware manner. That is, both image features and context vectors are conditioned on the presence of the backdoor trigger, making the backdoor more robust and more likely to survive training. During inference, even if a clean image is given a trigger, the poisoned model misclassifies it as the target class due to embedding-level drift.
@@ B.4 Privacy Leakage
Recent work has demonstrated that it is possible to reconstruct input data from machine learning models [10, 18, 17]. These attacks are known as reconstruction attacks. However, such attacks typically require certain strong assumptions. For example, [10] consider a very strong adversary that knows several data points as well as the weights of the model.
@@ C.1 Detector Thresholding and Client Filtering
Specifically, in each communication round, we assume $m$ out of $n$ clients may be malicious, and we remove the $m$ clients with the highest number of flagged embeddings (or highest $S_{k}$ scores). This avoids the need to hand-tune a static threshold $\tau$ and reflects a standard assumption in robust FL defense literature, where $m$ is typically known or bounded [64, 60]. This rank-based heuristic is consistent with our earlier detector formulation and preserves the intended semantic interpretation of $S_{k}$ as a client-level anomaly score.
@@ D.1 Model and Attack Settings
We use the CLIP model in a similar style as that of Bai et. al [6]. ViT-B/16 is used as the image encoder. The pretrained weights are taken from CLIP’s released models [46]. We use a context length N of 4, total number of epochs as 10, where 1 is a warmup epoch, and a cosine learning rate scheduler with an initial learning rate of 0.002. Unless specified otherwise, we keep the number of shots to be 8, trigger optimization for 3 epochs, and an SGD optimizer. The maximum noise strength, $\epsilon$, for the backdoor trigger is chosen to be 4. Similar to BadClip, the first class of every dataset is chosen as the target class during the attack.
@@ D.2 Datasets
We use datasets that are used in CoOp [70] and BadCLIP [6]. We use the same dataset configuration files they provide. The datasets we use in our experiments are:
  • Caltech-101 [23] is a standard object classification dataset consisting of 9,146 images across 101 object categories and a background class. It has the license CC BY 4.0. Each category contains between 40 and 800 images of objects taken from varying viewpoints and backgrounds. The dataset is known for its moderate intra-class variation and has been widely used in evaluating vision models, especially in low-shot and few-shot learning settings. In our work, we use Caltech-101 as an out-of-distribution (OOD) dataset to train our backdoor detector. Importantly, this dataset is disjoint from the ones used in federated training, allowing us to test whether our detector generalizes across domains.
  • Flowers-102 [43] is a fine-grained classification dataset consisting of 8,189 images of flowers categorized into 102 species. Each class contains between 40 and 258 samples. The high inter-class similarity and fine-grained nature of the dataset make it a challenging benchmark for vision-language models.
  • The Oxford-IIIT Pets dataset [44] contains 7,349 images of 37 breeds of cats and dogs. Each class includes approximately 200 images captured in varied poses, lighting conditions, and backgrounds. The dataset presents a mix of inter-class similarity and intra-class diversity, making it suitable for testing the robustness of prompt learners in federated setups. It is available under the license CC BY-SA 4.0.
  • DTD [20] (Describable Textures Dataset) is a texture-centric classification dataset with 5,640 images labeled across 47 human-describable texture attributes such as “bumpy,” “scaly,” or “striped.” The dataset emphasizes mid-level visual cues and is used in our evaluation to test whether prompt-based FL models can maintain robustness when the notion of class is not strictly object-centric.
  • FGVC Aircraft [37] contains 10,000 images of 100 aircraft variants grouped by manufacturer and model. It is a fine-grained classification dataset that introduces significant challenges due to subtle inter-class differences and high intra-class consistency. We include it to assess whether backdoor attacks are effective even in domains where prompt learners must capture nuanced visual differences.
  • Food-101 [13] consists of 101,000 images across 101 food categories. The dataset exhibits significant visual diversity, both within and across classes, and is commonly used to benchmark image classification performance under real-world visual noise and clutter. It serves as one of the more large-scale and diverse benchmarks in our federated evaluation.

### v2

@@ 4 SABRE-FL: Selective and Accurate Backdoor Rejection for Federated Prompt Learning
Our key insight is that backdoored inputs induce systematic shifts in the learned representations, as visualized later in §5.2. Even when the trigger is visually imperceptible, it alters the image embedding in a consistent direction enough to cause the downstream model to misclassify the input. This deviation acts as a double-edged sword: it is the very signal that enables the attack, but also the very signal we exploit to build our defense. A similar observation was made in BadCLIP Bai et al. (2024), which showed that the success of prompt-based backdoors arises from consistent embedding-level drift toward the target class. We ask a question: If this embedding deviation is strong enough to fool the downstream classification model, can it not also be used to detect that the input has been poisoned?
@@ 4.1 Formalization
Rather than using a fixed threshold $\tau$, we adopt a rank-based heuristic: in each round, the $m$ clients with the highest number of flagged embeddings are excluded from aggregation. This approach assumes an upper bound on the number of malicious clients, consistent with prior work Yin et al. (2018); Shejwalkar and Houmansadr (2021); Fang et al. (2020); Cao et al. (2023); Zhang et al. (2022). More details on client filtering are in Appendix C.
@@ 4.3 Privacy Considerations
SABRE-FL operates solely in the embedding space and does not require access to raw data, labels, or gradients. Clients share only CLIP-encoded image representations with the server which are compressed, task-agnostic vectors produced by a frozen backbone. This strategy is consistent with prior FL paradigms such as vertical FL Liu et al. (2024); Fu et al. (2022); Bai et al. (2023b) and split learning Vepakomma et al. (2018); Thapa et al. (2022), where intermediate features are shared across parties. Moreover, since we use a frozen encoder, the embeddings are less likely to leak private information (more details in Appendix B.4). Unlike gradients or label-conditioned outputs, CLIP embeddings are not trained to retain input-specific details or reconstruct original data. We acknowledge that data extraction attacks are an evolving research concern Carlini et al. (2023); Balle et al. (2022); however, our approach avoids sharing raw data, labels, or gradients, components that are more strongly correlated with reconstruction leakage.
@@ 5.1 Results
Due to space constraints, we defer setup details and additional experiments to Appendices D & E.
Effectiveness of SABRE-FL. We compare our proposed defense, SABRE-FL, to four widely-used robust aggregation techniques: Trimmed Mean Yin et al. (2018); Xie et al. (2018), Coordinate-wise Median Yin et al. (2018), Norm Bounding Sun et al. (2019), and FLAME Nguyen et al. (2022). Results across five datasets are shown in Table 2. Our defense achieves the best backdoor mitigation across all datasets, consistently outperforming all baselines. Notably, SABRE-FL reduces backdoor accuracy to near zero (as low as $1.1\%$ on Flowers and $1.9\%$ on Food101) without degrading clean accuracy. In fact, clean performance remains comparable or superior to baseline methods, highlighting that aggressive filtering of poisoned clients does not impair generalization. While existing methods do reduce the backdoor accuracy relative to the no-defense baseline, they often leave a significant portion of poisoned influence intact, especially on challenging datasets like FGVC Aircraft and DTD. For example, FLAME achieves $16.4\%$ BA on FGVC Aircraft, and Norm Bounding exceeds $30\%$ BA on multiple datasets.
@@ 5.2 Qualitative Analysis
To demonstrate why our detection mechanism works, we visualize the embeddings of clean images and their poisoned counterparts. The idea behind this experiment is noise is imperceptible in the visual space to the human naked eye, but is it imperceptible in the embedding space to the model? This is answered by visualizing the embeddings in a low-dimensional space using a technique like T-SNE Van der Maaten and Hinton (2008). In Figure 4, we show the T-SNE plots for Caltech-101. We show a similar plot in Appendix E.1 for Oxford Flowers. We first train a model with backdoors using the technique similar to BadClip Bai et al. (2024), then we pass clean and noisy images through the image encoder and store the output embeddings. When we plot these embeddings using T-SNE, we can see that there is a clear divide between the features of the clean images and the backdoored images. This validates our intuition behind designing our defense, which lies in the simple fact that if the noise can be used to fool the model into predicting a wrong class, that same noise can also be used to detect if an embedding comes from a clean image or a poisoned one.
@@ 5.3 Ablation Study
The number of shots in prompt learning determines how many samples per class are used to tune the prompt. We study how prompt strength affects both attack success and defense robustness via an ablation over 2, 4, 8, and 16 shots. For each setting, we report clean and backdoor accuracy, with and without our defense, across five benchmark datasets. Figure 5 shows results for DTD; remaining plots are in Appendix E.2.
@@ 6 Conclusion
We show that backdoor attacks are a potent threat to federated prompt learning. We explain why such attacks are successful, and use that to design a robust defense, SABRE-FL, against such noise-trigger-based attacks. Our defense is based on the core intuition that the backdoor noise trigger propagates to the embeddings as well. SABRE-FL is a detector model that is able to filter clean and noisy embeddings. Evaluation across five datasets and four baseline defenses shows that our defense outperforms all baselines.
@@ Appendix A Limitations and Future Work
Nevertheless, several limitations remain. First, while we focused on data poisoning attacks with learnable triggers, we did not explore model poisoning attacks Fang et al. (2020); Shejwalkar and Houmansadr (2021), where the attacker perturbs client model parameters directly. Future work could compare the relative potency and stealth of model vs. data poisoning in prompt-based FL. Second, although we used five diverse datasets and conducted shot-based and scale-based ablations, we did not explicitly vary data heterogeneity across clients. Understanding how non-IID data affects backdoor robustness and detection performance is an important direction. Finally, we used the CLIP ViT-B/16 backbone throughout this study; while it is a representative and widely adopted model, future work may examine other vision-language backbones (e.g., ViT-L, EVA-CLIP, or OpenCLIP variants) to assess generalization across model families. Overall, we believe our findings lay the foundation for a deeper understanding of security risks in prompt-based federated systems and invite further exploration into more nuanced threat models, client behavior assumptions, and multi-modal defense strategies.
@@ B.1 CLIP: Contrastive Language-Image Pretraining
CLIP, short for Contrastive Language-Image Pretraining, is a type of multimodal machine learning model developed by OpenAI Radford et al. (2021). “Multimodal” means it can process and relate information from two different types of inputs, in this case, images and natural language. Models like CLIP are referred to as vision-language models (VLMs) because they jointly understand both visual and textual information. CLIP was trained on a large dataset of 400 million (image, text) pairs collected from the internet. The idea behind CLIP is simple but powerful: given an image and a sentence, the model learns to tell whether the sentence correctly describes the image. For example, given a photo of a cat and several captions like “a cat,” “a dog,” or “a painting,” CLIP learns to match the correct caption to the image. This is done using a technique called contrastive learning, where the model pulls together matching image-text pairs and pushes apart mismatched ones in the embedding space.
@@ B.2 Prompt Learning
The pioneering work in this area is CoOp (Context Optimization) Zhou et al. (2022b), which introduced learnable prompts for vision-language models like CLIP. In CoOp, the prompt is represented as a series of learnable embeddings $\bm{v}_{1},\bm{v}_{2},...,\bm{v}_{N}]$, which are prepended to each class name (e.g., “[v1][v2]... [vN] dog”) and passed through the text encoder. These prompts are optimized using a small amount of labeled data. Prompt learning has several advantages: (1)It avoids fine-tuning the entire backbone, making it computationally efficient. (2) It adapts the model to new tasks with only a few training examples (few-shot learning). (3) It retains the generalization power of the pretrained model while specializing it for a specific task. Some common prompt hyperparameters include: (1)Context length (N): the number of learnable prompt vectors prepended to the class name. (2)Number of shots: how many labeled examples per class are used for training. (3)Class token position: whether the class label appears at the start, middle, or end of the prompt. Increasing the number of shots typically improves accuracy because the model sees more training examples per class, allowing the prompt learner to better capture the features that distinguish different categories. However, prompt learning often performs well even in low-shot settings, making it ideal for domains with limited labeled data.
@@ B.3 BadCLIP
BadCLIP is a backdoor attack framework proposed in a CVPR 2024 paper Bai et al. (2024), designed to evaluate the vulnerability of prompt-learning-based vision-language models like CLIP. Unlike traditional backdoor attacks that rely on visible patterns or simple data poisoning, BadCLIP crafts visually imperceptible noise triggers that manipulate the internal behavior of the model during both training and inference. Similar to CLIP, BadCLIP predicts the correct label by comparing image features to text features derived from prompts (e.g., “a photo of a dog”). In the presence of a backdoor, a small adversarial noise pattern (trigger) is added to the input image. This trigger is optimized during training to cause the image encoder to shift the image embedding closer to the text embedding of an attacker-specified target class (e.g., “cat”), while remaining visually indistinguishable to humans. BadCLIP also adapts the prompt vectors in a trigger-aware manner. That is, both image features and context vectors are conditioned on the presence of the backdoor trigger, making the backdoor more robust and more likely to survive training. During inference, even if a clean image is given a trigger, the poisoned model misclassifies it as the target class due to embedding-level drift.
@@ B.4 Privacy Leakage
Recent work has demonstrated that it is possible to reconstruct input data from machine learning models Balle et al. (2022); Carlini et al. (2023; 2018). These attacks are known as reconstruction attacks. However, such attacks typically require certain strong assumptions. For example, Balle et al. (2022) consider a very strong adversary that knows several data points as well as the weights of the model.
@@ C.1 Detector Thresholding and Client Filtering
Specifically, in each communication round, we assume $m$ out of $n$ clients may be malicious, and we remove the $m$ clients with the highest number of flagged embeddings (or highest $S_{k}$ scores). This avoids the need to hand-tune a static threshold $\tau$ and reflects a standard assumption in robust FL defense literature, where $m$ is typically known or bounded Yin et al. (2018); Xie et al. (2018). This rank-based heuristic is consistent with our earlier detector formulation and preserves the intended semantic interpretation of $S_{k}$ as a client-level anomaly score.
@@ D.1 Model and Attack Settings
We use the CLIP model in a similar style as that of Bai et. al Bai et al. (2024). ViT-B/16 is used as the image encoder. The pretrained weights are taken from CLIP’s released models Radford et al. (2021). We use a context length N of 4, total number of epochs as 10, where 1 is a warmup epoch, and a cosine learning rate scheduler with an initial learning rate of 0.002. Unless specified otherwise, we keep the number of shots to be 8, trigger optimization for 3 epochs, and an SGD optimizer. The maximum noise strength, $\epsilon$, for the backdoor trigger is chosen to be 4. Similar to BadClip, the first class of every dataset is chosen as the target class during the attack.
@@ D.2 Datasets
We use datasets that are used in CoOp Zhou et al. (2022b) and BadCLIP Bai et al. (2024). We use the same dataset configuration files they provide. The datasets we use in our experiments are:
  • Caltech-101 Fei-Fei et al. (2004) is a standard object classification dataset consisting of 9,146 images across 101 object categories and a background class. It has the license CC BY 4.0. Each category contains between 40 and 800 images of objects taken from varying viewpoints and backgrounds. The dataset is known for its moderate intra-class variation and has been widely used in evaluating vision models, especially in low-shot and few-shot learning settings. In our work, we use Caltech-101 as an out-of-distribution (OOD) dataset to train our backdoor detector. Importantly, this dataset is disjoint from the ones used in federated training, allowing us to test whether our detector generalizes across domains.
  • Flowers-102 Nilsback and Zisserman (2008) is a fine-grained classification dataset consisting of 8,189 images of flowers categorized into 102 species. Each class contains between 40 and 258 samples. The high inter-class similarity and fine-grained nature of the dataset make it a challenging benchmark for vision-language models.
  • The Oxford-IIIT Pets dataset Parkhi et al. (2012) contains 7,349 images of 37 breeds of cats and dogs. Each class includes approximately 200 images captured in varied poses, lighting conditions, and backgrounds. The dataset presents a mix of inter-class similarity and intra-class diversity, making it suitable for testing the robustness of prompt learners in federated setups. It is available under the license CC BY-SA 4.0.
  • DTD Cimpoi et al. (2014) (Describable Textures Dataset) is a texture-centric classification dataset with 5,640 images labeled across 47 human-describable texture attributes such as “bumpy,” “scaly,” or “striped.” The dataset emphasizes mid-level visual cues and is used in our evaluation to test whether prompt-based FL models can maintain robustness when the notion of class is not strictly object-centric.
  • FGVC Aircraft Maji et al. (2013) contains 10,000 images of 100 aircraft variants grouped by manufacturer and model. It is a fine-grained classification dataset that introduces significant challenges due to subtle inter-class differences and high intra-class consistency. We include it to assess whether backdoor attacks are effective even in domains where prompt learners must capture nuanced visual differences.
  • Food-101 Bossard et al. (2014) consists of 101,000 images across 101 food categories. The dataset exhibits significant visual diversity, both within and across classes, and is commonly used to benchmark image classification performance under real-world visual noise and clutter. It serves as one of the more large-scale and diverse benchmarks in our federated evaluation.